# Drillboards: Adaptive Visualization Dashboards for Dynamic Personalization of Visualization Experiences

### v1

@@ 5.2 Procedure
Phase II: Data Preparation. In this phase, the research team adapted the datasets provided by the experts so that they could be loaded into the DrillVis tool. While DrillVis nominally accepts any CSV file, there typically was some formatting and curation in this phase. For example, one participant provided an Excel spreadsheet with multiple worksheets; we split these into separate CSV files to use in the DrillVis tool. This phase lasted 3-4 days on average.
Phase III: Drillboard Authoring. When we met with our domain experts, we first demonstrated the authoring mode in the DrillVis tool using the farm dataset that they had seen in Phase I. During this demonstration, we showed all the aggregation and layout operators. We then let the experts themselves try out the tool and ask questions. Once they indicated they were ready to proceed, we provided them with a link to the DrillVis website with their dataset preloaded. We now asked the experts to use the authoring mode to create a new drillboard for their own data. We instructed them to author the drillboard so that it had two levels of detail that would be suitable for an expert and for a complete novice. We also asked them to list three in-depth questions about the dataset that should be answerable using the drillboard. Again, the experts followed a think-aloud protocol. There was no time limit during this phase, and we did not collect any metrics beyond observational notes and recordings. Participants were allowed to ask questions freely about the DrillVis tool. These sessions lasted between 1 hour to 1h and 30 minutes, and ended when the expert was satisfied with the product.
@@ 5.3.1 Authored Drillboards
Figure 12: P1’s drillboard P1 is analyzing financial assistance to a nation using two data sources. P1’s representation, shown above, is tailored for novices with greater abstraction for easier understanding.
@@ 5.3.2 Domain Experts
Unlike the OECD dataset, which covers country-based funding, the OCHA dataset targets financial inputs from global entities, including the UN and NGOs like Oxfam. P1 arranged the OCHA visualizations into groups reflecting his interests. Once the setup was finished, he scrutinized the differences between the two datasets.
@@ 6 Discussion and Limitations
Here we discuss the outcomes of our study, generalizing the insights gleaned from the deployment of DrillVis and examining its implications for the broader field of adaptive visualization dashboards. Our analysis reveals several key findings regarding user interaction with hierarchical data representations, personalization of data exploration, and the overall utility of drillboards in accommodating diverse user needs. We also identify limitations inherent to our current implementation and propose directions for future research.
@@ 6.1 Explaining the Results
Observations from experts. We observe that each expert aggregated their hierarchy with different intent. P1 used the tool for macro-level trend analysis by aggregating many charts for broader insights. P2 focused on using arithmetic operations to quickly detect issues. P3 highlighted differences between wealthy and impoverished regions by emphasizing a key feature. These findings suggest that such approach to progressive abstraction actually mirrors many expert’s mental models of such data; humans have a tendency to summarize and abstract, and drillboards support this model.
@@
6.2 Limitations and Future Work
End-user challenges with perceiving and navigating the drillboard may be explained by the lack of visual stability. The dynamic expansion and collapse of charts during drill-down and roll-up operations sometimes led to confusion, likely because the changing order and appearance of charts disrupted the participants’ mental map of the data. After trying three different versions of drillboards, two participants, E2, and E7, admitted confusion in when drilling down into various types of aggregate operators. E2 said, ‘The drill-down interaction is occasionally unpredictable; at times, it displays the data elements, sometimes decomposing a multi-line chart into individual lines, and at other times, presenting arithmetic operation elements. This lack of consistency leads to confusion.’
@@ 6.2 Limitations and Future Work
We also identify several methods to improve drillboards. One direction is to make them easier to use for a wider audience by providing effective guidance or automating the aggregation process. Furthermore, there are various charts that drillboards do not yet provide and numerous aggregation mechanisms that are not implemented. All these issues are left for future work, however.

### v3

@@ 5.2 Procedure
Phase II: Data Preparation. In this phase, the research team adapted the datasets provided by the experts so that they could be loaded into the DrillVis tool. While DrillVis nominally accepts any CSV file, there typically was some formatting and curation in this phase. For example, one participant provided an Excel spreadsheet with multiple worksheets; we split these into separate CSV files to use in the DrillVis tool. This phase lasted 3-4 days on average. The intent of giving 3 to 4 days was because these domain experts were extremely busy. It is not because it took 3 to 4 days, but we wanted them to do the study whenever they had time. Furthermore, the goal was to give them time to think of the data before designing their own personalized structure.
Phase III: Drillboard Authoring. When we met with our domain experts, we first demonstrated the authoring mode in the DrillVis tool using the farm dataset that they had seen in Phase I. During this demonstration, we showed all the aggregation and layout operators. We then let the experts themselves try out the tool and ask questions. Once they indicated they were ready to proceed, we provided them with a link to the DrillVis website with their dataset preloaded. We now asked the experts to use the authoring mode to create a new drillboard for their own data. We instructed them to author the drillboard so that it had two levels of detail that would be suitable for an expert and for a complete novice. We also asked them to list three in-depth questions about the dataset that should be answerable using the drillboard. Again, the experts followed a think-aloud protocol. There was no time limit during this phase, and we did not collect any metrics beyond observational notes and recordings. Participants were allowed to ask questions freely about the DrillVis tool. These sessions lasted between 1 hour to 1h and 30 minutes, and ended when the expert was satisfied.
@@
5.3 Usability Study
Figure 12: Drillboard usability study. We conducted usability tests using the PSSUQ and NASA-TLX questionnaires to evaluate user satisfaction and workload in Drillvis. The Mean indicates the arithmetic mean scores of all participants, and SD refers to their standard deviation. Three PSSUQ questions (7, 8, and 9) were excluded as they were not relevant to our tool. Results indicate that participants were generally satisfied, and NASA-TLX findings suggest the workload was manageable, though some found the task slightly mentally demanding.
As a post-study, we conducted a usability evaluation with participants using the PSSUQ and NASA-TLX questionnaires to assess satisfaction and workload in Drillvis. Figure 12 presents the results of this usability study. We excluded three PSSUQ questions that were not relevant to our system’s conditions: e.g., ‘7. The system gave error messages that clearly told me how to fix problems,’ ‘8. Whenever I made a mistake using the system, I could recover easily and quickly’), and ‘9. The information (such as online help, on-screen messages, and other documentation) provided with this system was clear.
Overall, participants were satisfied with using Drillvis, as reflected in the PSSUQ results. Most questions had an average score between 1 and 2, with little variation in standard deviation. The highest standard deviation was observed in Q15, which asked whether the system had all the expected functions and capabilities. Responses to this question varied slightly, though none of the participant expressed disagreement.
In the NASA-TLX questionnaire, participants generally did not find the workload overwhelming. However, responses regarding mental demand and effort were mixed, with some participants reporting a substantial cognitive load. This suggests that while the system was manageable for most, certain aspects required significant mental effort for some users. We believe this may be linked to the issue of visual stability. We further discuss this in Sec. 6.2.
@@ 5.4.1 Authored Drillboards
Figure 13: P1’s drillboard. P1 is analyzing financial assistance to a nation using two data sources. P1’s representation, shown above, is tailored for novices with greater abstraction for easier understanding.
@@ 5.4.2 Domain Experts
Unlike the OECD dataset, which covers country-based funding, the OCHA dataset targets financial inputs from global entities, including the UN and NGOs like Oxfam. P1 arranged the OCHA visualizations into groups reflecting his interests. Once the setup was finished, he scrutinized dataset differences.
@@ 6 Discussion and Limitations
Here we discuss the outcomes of our study, generalizing the insights gleaned from the deployment of DrillVis and examining its implications for the broader field of adaptive visualization dashboards. Our analysis reveals key findings regarding user interaction with hierarchical data representations, personalization of data exploration, and the overall utility of drillboards in accommodating diverse user needs. We also identify limitations inherent to our work and propose directions for future research.
@@ 6.1 Explaining the Results
Observations from experts. Each expert aggregated their hierarchy with different intent. P1 used the tool for macro-level trend analysis by aggregating many charts for broader insights. P2 focused on using arithmetic operations to quickly detect issues. P3 highlighted differences between wealthy and impoverished regions by emphasizing a key feature. This suggests that progressive abstraction actually mirrors many expert’s mental models of such data; humans have a tendency to summarize and abstract, and drillboards support this model.
@@
6.2 Limitations
Visual stability. End-user challenges with perceiving and navigating the drillboard may be explained by the lack of visual stability. The dynamic expansion and collapse of charts during drill-down and roll-up operations sometimes led to confusion, likely because the changing order and appearance of charts disrupted the participants’ mental map of the data. After trying three different versions of drillboards, two participants, E2, and E7, admitted confusion in when drilling down into various types of aggregate operators. E2 said, ‘The drill-down interaction is occasionally unpredictable; at times, it displays the data elements, sometimes decomposing a multi-line chart into individual lines, and at other times, presenting arithmetic operation elements. This lack of consistency leads to confusion.’
@@ 6.2 Limitations
Drillboards with different charts. While our approach is adaptable to different chart types, each type has unique aggregation conditions and functionalities that require tailored policies. For instance, a treemap—which is inherently hierarchical—could be constructed using Drillboards, but the way it summarizes data (e.g., grouping levels) differs from how line or bar charts aggregate information. Certain interactions, such as overlay or projection, may also be less applicable. Similarly, in node-link diagrams, operations like labeling, summarization, and archetyping seem feasible, but projection and overlay may not work as intuitively. These differences indicate that Drillboards’ aggregation model is not universally applicable without modification. Therefore, while Drillboards can support various visualizations, each implementation requires specific adaptations, which we leave as future work.
6.3 Future Work
@@
We also identify methods to improve drillboards. One direction is to make them easier to use for a wider audience by providing effective guidance or automating the aggregation process. Furthermore, there are various charts that drillboards do not yet provide and numerous aggregation mechanisms that are not implemented. All these issues are left for future work, however.